Circular No. 8516

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005br IN IC 5084 Further to IAUC 8515, L. A. G. Monard reports his discovery of an

apparent supernova (mag ∼ 16.5 ± 0.3) on unfiltered CCD images taken on Mar. 28.113 UT and confirmed on Apr. 19.114 (at mag ∼ 16.4 ± 0.2). The new object is located at α = 21^{h}09^{m}13^{s}.09, δ = −63^{o}17′05′′.6 (equinox 2000.0), which is 14′′ west and 15′′ north of the nucleus of IC 5084. SN 2005br is not visible on an image from the Digitized Sky Survey (limiting red mag 20.5) or on an image taken by Monard on Mar. 5.132 (limiting red mag 18.0).

COMETS C/2005 D2–D5, E5–E9, F1, F2 (SOHO) Following are the initial positions for additional Kreutz sungrazing

comets found on SOHO website images (cf. IAUC 8515). K. Battams reports that C/2005 D5 reached peak brightness near mag 5–6; the other Feb. objects were only a little above background (peak mag 7–8). C/2005 D2 was stellar and faded out rapidly at 8–9R_{⊙}; C/2005 D3 was brighter and also stellar, fading rapidly at ∼ 9R_{⊙}; C/2005 D4 was small and faded at 9– 10R_{⊙}. No tail was visible for C/2005 D2, D3, D4, E7, and F2. Tail lengths (uncertainties ± 1′) for C/2005 D5: tail developed at ∼ 12R_{⊙}; at 6.3R_{⊙}, > 10′ (head very much smaller); at 5.9R_{⊙}, 16′.8 (head effectively gone); at 5.7R_{⊙}, 16′ (no head). Physical information for other comets: C/2005 E6, diffuse, possible tail; C/2005 E7, very faint, diffuse; C/2005 E8, diffuse with a faint tail 262′′ long at ∼ 7.3R_{⊙}; C/2005 E9, very small and faint; C/2005 F1, faint and diffuse, with hint of a tail; C/2005 F2, stellar.

Comet  2005 UT  α_{2000}  δ_{2000} Inst. F  MPEC
C/2005 D2 Feb. 22.087 22^{h}37.^{m}6 −11^{o}01′ C3  TH 2005-G40
C/2005 D3  22.279 22 44.9 −11 31 C3  KC 2005-G40
C/2005 D4  26.346 22 55.1 − 9 34 C3  BZ 2005-G40
C/2005 D5  26.571 23 07.9 −10 18 C3/2 BZ 2005-G40
C/2005 E5 Mar. 11.966 23 45.9 − 4 14 C3  HS 2005-G93
C/2005 E6  13.929 0 03.1 − 4 03 C3/2 BZ 2005-G93
C/2005 E7  14.404 23 50.4 − 3 41 C3/2 XL 2005-G93
C/2005 E8  14.863 0 01.4 − 3 21 C3/2 TH 2005-G93
C/2005 E9 14.992 23 47.4 − 2 36 C2 BZ 2005-G93
C/2005 F1  17.029 0 12.3 − 2 43 C3/2 HS 2005-G94
C/2005 F2  18.571 0 10.8 − 1 44 C3/2 HS 2005-G94

2005 April 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green